OpenStudy (iluvsoccer):

what is the solution to the system of equations? 5x+4y=8 x-2y=10

OpenStudy (anonymous):

multiply the second equation all the way across by 2, then add it to the first one

OpenStudy (triciaal):

to use the addition/elimination method choose one of the variables and use the appropriate factors to make the coefficients equal and opposite so that when you add the equations the variable is eliminated. you have one linear equation with one unknown. solve. use this value in either of the originals to get the value of the other variable.

OpenStudy (triciaal):

as suggested by @satellite73 multiply the 2nd equation by 2 to eliminate y's when added to the 1st equation

OpenStudy (anonymous):

5x+4y=8 a) x-2y=10 b) multiplying b by 2 will result in 2x-4y=20 c) add c and a 2x-4y=20 5x+4y=8 7x=28 x=4 put x=4 in b) 4-2y=10 -2y=6 y=-3 x=4 , y=-3